OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP ME WITH THIS WORD PROBLEM! IDK WHAT TO DO AT ALL! The number of milligrams of a drug that remains in a patient's system after t hours is given by the function A(t) = Iert. Juan was given 500 milligrams of medicine which leaves his bloodstream at a rate of 20%. How much of the medicine remains in his system after 6 hours? 150.60 mg 133.90 mg 111.58 mg 99.20 mg

OpenStudy (aaronq):

you can do this by simply subtracting 20 percent each time, it's similar to half-lifes (if you've ever done that) hour 0: 200*0.2=40 hour 1: 160*0.2=32 hour 2: 160-32=128 etc

OpenStudy (aaronq):

sorry change the 200 to 500

OpenStudy (aaronq):

hour 0: 500 hour 1: 500*0.2=100 -> 500-100=400 left hour 2: 400*0.2=80 -> 400-80=320 left hour 3: 320*0.2=64 -> 320-64= 256 left

OpenStudy (anonymous):

I am still lost..

OpenStudy (whpalmer4):

It might be easier if you think of this in terms of what is left, rather than what is leaving. If 20% leaves each hour, then after each hour passes, you have 100%-20% = 80% of the previous amount remaining. 80% = 80/100 = 0.8 At hour 0, you have 500 At hour 1 (1 hour elapsed), you have 500*0.8 = 400 At hour 2 (2 hours elapsed) you have 400*0.8 = 320 At hour 3 (3 hours elapsed) you have 320*0.8 = 256 and so on. Notice that you can also write this progression as 500 = 500 500*0.8 = 400 500*0.8*0.8 = 320 500*0.8*0.8*0.8 = 256 etc. which is what the formula you were given does.

OpenStudy (whpalmer4):

Are you familiar with compound interest? It's the same underlying concept.

OpenStudy (anonymous):

not really..

OpenStudy (anonymous):

is it gonna be b?

OpenStudy (whpalmer4):

Get out your calculator. multiply 0.8*0.8*0.8*0.8*0.8*0.8*500. That's the amount left. Someone needs to calculate it to get the answer, it might as well be you :-)

OpenStudy (anonymous):

I did the closest I got was to B

OpenStudy (whpalmer4):

That must be it then!

OpenStudy (anonymous):

its 121.072. kay thank ya!

OpenStudy (whpalmer4):

131.072, actually